# Breaking the Reviewer: Assessing the Vulnerability of Large Language Models in Automated Peer Review Under Textual Adversarial Attacks

### v1

@@ Appendix G Evaluation – Adversarial Attack Result for PeerRead Dataset
Table 6: Adversarial Attack Performance and Score Shift Comparison of PeerRead Kang et al. (2018) dataset. “ASR” represents the Attack Success Rate. “Avg. Score shift” signifies the average score shift between attacked outputs and clean outputs. “Avg. #Pos/Neg Shift” indicates the average change in the number of positive and negative tags assigned to the reviews. “#Queries” refers to the average number of querying LLM. The value are reported as result from GPT-4o-mini / GPT-4o/ Llama-3.3-70B / Mistral-Small-3.1.
Attacks ASR $\uparrow$ Average Shift #Queries
@@
DeepWordBug 0.92 / 0.81 / 0.82 / 0.89 4.27 / 3.87 / 3.06 / 3.22 0.33 / 0.00 / 0.07 / -0.02 -0.95 / 1.18 / -1.73 / -0.24 42 / 51 / 50 / 56
PuncAttack 0.87 / 0.90 / 0.79 / 0.83 3.90 / 4.23 / 3.17 / 2.98 0.33 / -0.03 / 0.03 / -0.02 -0.81 / 1.18 / -1.71 / -0.20 51 / 63 / 64 / 68
TextFooler 0.92 / 0.95 / 0.87 / 0.91 4.61 / 4.30 / 2.87 / 3.44 0.35 / -0.03 / -0.07 / -0.01 -0.90 / 1.50 / -2.14 / -0.32 46 / 54 / 57 / 58
BERT-Attack 0.85 / 0.74 / 0.70 / 0.85 4.13 / 4.43 / 2.35 / 3.64 0.28 / 0.03/ -0.09 / -0.02 -0.75 / 1.23 / -1.70 / -0.40 52 / 53 / 54 / 65
StyleAdv 0.83 / 0.88 / 0.80 / 0.87 3.58 / 4.28 / 2.35 / 2.81 0.27 / -0.10 / -0.14 / -0.02 -0.65 / 1.13 / -1.96 / -0.26 12 / 16 / 12 / 16
Table 6 presents the results of different attack methods aimed at intentionally boosting the scores assigned by the LLM-based reviewer. Based on the results, we draw the same trend as shown in Table 2. This demonstrates the effectiveness of the attack across different datasets and further confirms the vulnerability of LLMs to simple textual adversarial attacks under a black-box scenario.
A closer analysis of the results reveals that:
(1) TextFooler consistently achieves the highest Attack Success Rate (ASR) across models, with a peak of 0.95 for GPT-4o, suggesting that even minor lexical substitutions can significantly distort LLMs’ judgment. This high ASR is paired with a substantial average score shift, particularly for GPT-4o-mini, indicating that such perturbations can systematically induce inflated evaluations. The consistent increase in positive tags and reduction in negative tags further highlights a dual manipulation effect—enhancing positive sentiment while suppressing critical feedback—directly compromising review integrity, aligning with the increased predicted score.
@@
(2) StyleAdv, while maintaining a moderate ASR, achieves comparable score shifts with the lowest query count, making it the most efficient attack among the tested methods. This suggests that stylistic manipulations, even with minimal transformation candidates and model interaction, are particularly deceptive to LLMs, likely because of their characteristic in high modification rate and thus have space to alter the distribution of attacked content to differentiate from the original one.
@@
(3) GPT-4o and GPT-4o-mini exhibit the highest susceptibility to adversarial attacks, Llama-3.3-70B and Mistral-Small-3.1 show relatively better resilience. We hypothesize this difference arises from GPT models’ stronger emphasis on maintaining contextual coherence. This strength in natural text generation makes them more prone to retaining even sentiment shifts caused by manipulations as coherent opinions, leading to more pronounced score shifts. In contrast, Llama-3.3-70B’s architecture may prioritize semantic understanding over coherence, making it slightly more robust against such manipulations.
These findings emphasize the critical need for developing more robust LLMs, especially in high-stakes applications such as automated review generation. Beyond improving model architectures against existing textual adversarial attacks, future research should explore more resilient prompting techniques and processing mechanisms that can mitigate the impact of adversarial perturbations.
Appendix H Evaluation – Human Evaluation
@@
To answer these questions, we designed two complementary human studies: one focused on review quality evaluation, and another on attack stealthiness and identifiability. We collected user study results from 31 participants, all of whom are master’s or PhD students with background knowledge in deep learning and natural language processing. The setup and methodology is described below for each in turn.
@@ H.1 Review Quality Ranking
To evaluate human-perceived review quality, we adopt the following three criteria:
@@
(2) Content Relevance: Does the review engage with the paper’s core contributions, and does it possess alignment with the meta-review context?
@@
For each review set, participants were asked to rank the four anonymized reviews (one LLM-generated and three human-written) along the above three quality dimensions. Following each ranking task, they also provided a confidence score (on a 1–5 scale) indicating the certainties about their judgments.
@@
Result & Discussion. The human evaluation results in Table 7 reveal insights emerging regarding the human-perceived quality of LLM-generated reviews as follows:
(1) Lamas-3.3-70B consistently outperformed other models in all three ranking criteria, achieving the highest ratings (lowest Rank value) and the highest T-1 rate for Content Relevance, suggesting that it produces reviews most closely aligned with the core contributions of the evaluated articles. This superior performance aligns with the hypothesis that larger models are more proficient in grasping complex content.
(2) GPT-4o displayed a relatively weak performance in Insightfulness compared to other LLMs, with the lowest T-1 Rate and the highest average rank, indicating that it struggles to offer constructive, in-depth feedback, a critical factor in effective peer review. A potential explanation lies in GPT-4o’s strict adherence to our structured prompt, which encouraged balanced coverage of all review aspects but may have inadvertently constrained its ability to deliver focused, in-depth analysis of specific aspects. Notably, despite this limitation, GPT-4o’s reviews still outperformed at least one human-written review in the ranking tasks, highlighting its overall competitiveness.
@@
(3) Mistral-small-3.1, despite demonstrating strong Fluency & Consistency with the highest T-1 Rate, lagged in Insightfulness, suggesting that fluency alone does not guarantee meaningful critique.
@@
(4)GPT-4o-mini demonstrated a balanced performance, achieving competitive results across all criteria, especially in Insightfulness, indicating that smaller, more efficient models can still achieve strong human-perceived quality.
@@
These results collectively highlight the importance of balancing fluency, relevance, and depth in LLM-generated reviews, with Llama-3.3-70B setting a strong benchmark in this regard. However, the somewhat underwhelming performance of models like GPT-4o in providing insightful feedback underscore the ongoing challenge of enhancing LLMs to generate not just coherent text but also substantive and contextually valuable content.
@@
Table 7: User Study on Review Quality Ranking. “Rank” denotes the average position of the LLM-generated review among the three human reviews (lower is better), “T-1 Rate” is the proportion of times the LLM review was ranked first for that criterion. “Conf.” is the average confidence score users assigned to their rankings.
@@
H.2 Human-based Attack Imperceptibility Evaluation
To further examine the robustness of LLM-generated content, we conducted a human evaluation targeting the identifiability of adversarial perturbations under realistic review scenarios. Specifically, this experiment aimed to answer two key research questions:
(1) Attack Identification: In a simulated peer-review setting, can human evaluators distinguish which papers has been adversarially attacked?
@@ H.2 Human-based Attack Imperceptibility Evaluation
Settings. To this end, we recruited annotators with experience in academic writing to participate in a two-stage evaluation task. In the first stage, each annotator was presented with a pair of papers, one clean and one adversarially perturbed, and asked to identify which appeared adversarially manipulated. In the second stage, annotators were shown a single adversarially perturbed article and instructed to highlight the paragraph they believed was most indicative of adversarial editing via a given specific textual attack method.
@@
Result & Discussion. According to Table 8, we observe substantial variation in the imperceptibility of different granularities of adversarial attacks and discuss them as follows:
@@
(1) BERT-Attack emerges as the most stealthy method, with annotators achieving only 50% accuracy in identifying perturbed papers (i.e., chance level), and a low localization accuracy of 8%. This suggests that semantic-preserving synonym substitution strategies are particularly effective in preserving the naturalness of text under human scrutiny, making them difficult to detect even when readers are explicitly informed of adversarial presence.
(2) Despite being originally designed to subtly change writing style, StyleAdv showed a relatively high identification accuracy (73%) but notably poor localization performance (21%), accompanied by the highest uncertainty rate (34%). This outcome suggests that stylometric perturbations may not blend well in academic contexts, possibly because the injected styles (i.e., biblical) deviate too far from the expected scholarly tone.
@@
As detailed in Appendix F in our revision, StyleAdv also involves higher modification rates and lower semantic similarity, which may disrupt paragraph-level cohesion and topical consistency, making the manipulated version stand out when compared side-by-side with a clean reference. However, the poor localization accuracy indicates that in zero-shot settings, where no clean version is available for comparison, humans struggle to pinpoint the attack, highlighting the stealthy nature and potential risk of such stylistic perturbations. This threat may be further amplified if future attacks adopt writing styles that are more plausible within academic domains, such as those resembling non-native writing patterns.
@@
(3) PuncAttack resulted in high identification accuracy, which is expected—as even minimal edits like punctuation noise become salient when directly compared to a clean version, making extraneous or altered characters visually stand out. However, its moderate localization accuracy suggests that while readers often sense something is off, they struggle to pinpoint the exact manipulated paragraph—especially when such subtle noise is sparsely distributed across a lengthy article.
Besides, in real-world scenarios without prior knowledge of an attack, such perturbations may be easily dismissed as natural typos or stylistic quirks, allowing them to evade scrutiny. Thus, despite being simple, PuncAttack retains strong potential as a subtle and deceptive adversarial strategy.
(4) TextFooler and DeepWordBug both occupy a middle ground but differ in performance characteristics. DeepWordBug does not introduce obvious special characters like punctuation, making its subtle character-level perturbations difficult to identify during comparisons of lengthy articles, which results in relatively low identification accuracy.
@@
However, when annotators are informed about the presence of attacks and carefully examine the text, they achieve a high localization accuracy (96%), indicating that these spelling changes can be discovered with focused reading. In contrast, TextFooler exhibits lower identification and localization accuracies, suggesting that synonym substitution attacks are harder to detect due to their higher semantic preservation, which reduces their perceptibility.
@@
These findings highlight that the detectability of adversarial attacks varies significantly depending on their nature and granularity. Attacks that preserve semantics tend to be more stealthy, while those introducing stylistic or punctuation anomalies are easier to spot in direct comparisons but still pose risks in real-world, zero-knowledge scenarios. This underscores the ongoing challenge in designing robust defenses against diverse and evolving adversarial strategies.

### v3

@@ G.2 Evaluation – Attack Effectiveness
Table 7: Adversarial Attack Performance and Score Shift Comparison of PeerRead Kang et al. (2018) dataset. “ASR” represents the Attack Success Rate. “Avg. Score shift” signifies the average score shift between attacked outputs and clean outputs. “Avg. #Pos/Neg Shift” indicates the average change in the number of positive and negative tags assigned to the reviews. “#Queries” refers to the average number of times LLM is queried. The values are reported as result from GPT-4o-mini / GPT-4o/ Llama-3.3-70B / Mistral-small-3.1.
Attacks ASR $\uparrow$ Average Shift #Queries $\downarrow$
@@
DeepWordBug 0.92 / 0.81 / 0.82 / 0.89 4.27 / 3.87 / 3.06 / 3.22 0.33 / 0.00 / 0.07 / -0.02 -0.95 / -1.18 / -1.73 / -0.24 42 / 51 / 50 / 56
PuncAttack 0.87 / 0.90 / 0.79 / 0.83 3.90 / 4.23 / 3.17 / 2.98 0.33 / -0.03 / 0.03 / -0.02 -0.81 / -1.18 / -1.71 / -0.20 51 / 63 / 64 / 68
TextFooler 0.92 / 0.95 / 0.87 / 0.91 4.61 / 4.30 / 2.87 / 3.44 0.35 / -0.03 / -0.07 / -0.01 -0.90 / -1.50 / -2.14 / -0.32 46 / 54 / 57 / 58
BERT-Attack 0.85 / 0.74 / 0.70 / 0.85 4.13 / 4.43 / 2.35 / 3.64 0.28 / 0.03 / -0.09 / -0.02 -0.75 / -1.23 / -1.70 / -0.40 52 / 53 / 54 / 65
StyleAdv 0.83 / 0.88 / 0.80 / 0.87 3.58 / 4.28 / 2.35 / 2.81 0.27 / -0.10 / -0.14 / -0.02 -0.65 / -1.13 / -1.96 / -0.26 12 / 16 / 12 / 16
Table 7 shows the results of different attacks designed to boost scores assigned by the LLM-based reviewer. The trends align with those in Table 2, demonstrating the attacks’ effectiveness across datasets and the vulnerability of LLMs to simple textual adversarial attacks in a black-box setting. A closer analysis reveals that:
TextFooler effectively manipulates LLM evaluations across models. It consistently achieves the highest ASR, peaking at 0.95 for GPT-4o, showing that even minor lexical substitutions can significantly distort judgments. In general, positive tags tend to increase relative to negative tags, with the decrease in negative tags usually larger than the increase in positive tags. This biases reviews toward favorable evaluations and suppresses critical ones, undermining review integrity. The pattern does not hold for every model, as prompts that encourage both positive and negative feedback and emphasize critical evaluation can cause models that already assign high positive and low negative scores to reduce positive tags or increase negative tags after the attack, reflecting the combined effect of the prompt and the perturbation.
@@
StyleAdv is the most efficient attack among the tested methods. Despite achieving only moderate ASR, it produces noticeable score shifts with the lowest query count. This indicates that stylistic manipulations, even with minimal transformation candidates and limited model interaction, can effectively deceive LLMs. Their high modification rate allows the attacked content to differ sufficiently from the original, altering its distribution in a way that impacts evaluations.
@@
GPT models are susceptible to adversarial attacks, while Llama-3.3-70B shows better resilience. We hypothesize that this difference stems from GPT models’ stronger emphasis on maintaining contextual coherence. This strength in natural text generation makes GPT models more likely to interpret even manipulated sentiment shifts as coherent opinions, resulting in more pronounced score changes. In contrast, Llama-3.3-70B may prioritize semantic understanding over coherence, which provides slightly greater robustness against such manipulations.
Overall, these results underscore the importance of advancing LLM robustness and point to resilient prompting and processing strategies as promising directions for safeguarding review integrity against adversarial perturbations.
Appendix H Human Evaluation
@@
To answer these questions, we designed two complementary human studies: one focused on review quality evaluation, and another on attack stealthiness and identifiability. We collected user study results from 29 participants, all of whom are master’s or PhD students with background knowledge in deep learning and natural language processing. The setup and methodology for each in turn are described below.
@@ H.1 Review Quality Ranking
H.1.1 Evaluation Criteria
To assess review quality as perceived by humans, we employ the following three criteria:
@@ H.1.1 Evaluation Criteria
(2) Content Relevance: Does the review address the paper’s core contributions and align with the ground-truth meta-review context?
@@
H.1.2 Annotation Process.
To keep the task manageable and avoid annotator fatigue, we presented the meta-review of each paper, which served as a concise conference-style summary instead of the full manuscript. This meta-review provided the context for evaluating the relevance, coherence, and insightfulness of the reviews.
@@
Participants assessed bundles of four reviews, each consisting of one LLM-generated review (from GPT-4o-mini, GPT-4o, Llama-3.3-70B, or Mistral-small-3.1) and three human-written reviews. Every bundle was independently judged by all 29 participants. For each bundle, annotators ranked the four anonymized reviews along the three quality dimensions. After completing each ranking task, they also reported their confidence on a 1–5 scale, reflecting the certainty of their judgments.
H.1.3 Result & Discussion.
@@
The human evaluation results in Table 8 reveal insights emerging regarding the human-perceived quality of LLM-generated reviews as follows:
Llama-3.3-70B consistently excels in fluency and content relevance. It secured the highest rankings (lowest Rank values) in both Fluency & Consistency and Content Relevance. While its Insightfulness was slightly below that of GPT-4o-mini, it remained competitive and outperformed GPT-4o across most dimensions, highlighting its overall strength as a balanced review generator.
GPT-4o underperforms in insightfulness and overall constructive feedback. It showed weaker performance across all three criteria compared to other LLMs, with the lowest T-1 Rate and highest average Rank indicating difficulty in providing in-depth, constructive feedback. This limitation is not simply due to model size, as both GPT-4o-mini and Llama-3.3-70B performed better. The likely explanation is GPT-4o’s strict adherence to the structured prompt: while this ensured balanced coverage, it often produced exhaustive but surface-level comments, reducing sharp, focused critique. In contrast, GPT-4o-mini followed the prompt less rigidly, generating more targeted and insightful feedback, which human judges rewarded with higher T-1 Rates. Despite its shortcomings, GPT-4o still outperformed at least one human-written review in ranking tasks, demonstrating overall competitiveness.
@@
Mistral-small-3.1 excels in fluency but lacks insightfulness. Although it achieved the highest T-1 Rate for Fluency & Consistency, it lagged in Insightfulness, suggesting that fluency alone does not guarantee meaningful critique.
@@
GPT-4o-mini achieves balanced and insightful performance across criteria. It delivered competitive results in all evaluated dimensions, particularly in Insightfulness, indicating that smaller, more efficient models can still achieve strong human-perceived review quality.
@@
These results collectively highlight the importance of balancing fluency, relevance, and depth in LLM-generated reviews, with Llama-3.3-70B setting a strong benchmark in this regard. However, the somewhat underwhelming performance of models like GPT-4o in providing insightful feedback underscores the ongoing challenge of enhancing LLMs to generate not just coherent text but also substantive and contextually valuable content.
@@
Table 8: User Study on Review Quality Ranking. “Rank” denotes the average position of the LLM-generated review among the three human reviews (lower is better), and “T-1 Rate” is the proportion of times the LLM review was ranked first for that criterion. “Conf.” is the average confidence score users assigned to their rankings.
@@
H.2 Attack Imperceptibility Evaluation
To further examine the robustness of LLM-generated content, we conducted a human evaluation targeting the identifiability of adversarial perturbations under realistic review scenarios.
H.2.1 Evaluation Criteria.
Specifically, this experiment aimed to answer two key research questions:
(1) Attack Identification: In a simulated peer-review setting, can human evaluators distinguish which papers have been adversarially attacked?
@@ H.2 Attack Imperceptibility Evaluation
H.2.2 Annotation Process.
We recruited annotators with prior experience in academic writing to participate in a two-stage evaluation. In the first stage, annotators were presented with pairs of papers, one original and one adversarially perturbed, and were asked to determine which paper had been manipulated. In the second stage, annotators were provided with a single adversarially perturbed paper and were instructed to identify the paragraph they judged to be most indicative of adversarial modification under a specified textual attack method.
@@
H.2.3 Result & Discussion.
As shown in Table 9, the imperceptibility of adversarial attacks varies substantially across different granularities, as discussed below.
@@
BERT-Attack is the most stealthy adversarial method. Annotators achieved only 50% accuracy in identifying perturbed papers (i.e., chance level) and a low localization accuracy of 8%. This indicates that semantic-preserving synonym substitution strategies are highly effective in maintaining naturalness under human scrutiny, making adversarial changes difficult to detect even when readers are explicitly aware of their presence.
@@
StyleAdv is easier to identify but difficult to localize. Although it was designed to subtly alter writing style, it showed relatively high identification accuracy (73%) yet poor localization accuracy (21%), along with the highest uncertainty rate (34%). This suggests that stylometric perturbations do not blend well in academic contexts, likely because injected styles (e.g., biblical) deviate too far from scholarly tone. As detailed in Appendix C, StyleAdv also involves higher modification rates and lower semantic similarity, which may disrupt paragraph-level cohesion and topical consistency, making manipulations stand out when compared side-by-side with a clean version. However, the poor localization accuracy indicates that in zero-shot settings without a clean reference, humans struggle to pinpoint the attack, highlighting its stealthiness. This risk could grow if future attacks adopt writing styles more plausible in academic contexts, such as those resembling non-native writing patterns.
@@
PuncAttack leaves obvious patterns yet hinders accurate localization. It yielded high identification accuracy, which is expected since even minimal punctuation edits become salient when compared with a clean version. However, its moderate localization accuracy suggests that while readers sense something unusual, they struggle to pinpoint the exact manipulated paragraph, especially when subtle noise is sparsely distributed across a lengthy article. In real-world settings without prior knowledge of an attack, such perturbations may be dismissed as typos or stylistic quirks, allowing them to evade scrutiny. Despite its simplicity, PuncAttack retains strong potential as a subtle and deceptive adversarial strategy.
@@
TextFooler and DeepWordBug occupy the middle ground but differ in detectability. DeepWordBug avoids obvious cues such as punctuation, making its character-level perturbations hard to detect across lengthy paper content, which results in relatively low identification accuracy. Yet when annotators know attacks are present and examine carefully, localization accuracy rises to 96%, showing that spelling changes can be discovered with focused reading. In contrast, TextFooler exhibits lower identification and localization accuracies, suggesting that synonym substitution attacks are harder to detect due to higher semantic preservation, which reduces perceptibility.
@@
These findings show that attack detectability depends heavily on nature and granularity: semantic-preserving attacks are more stealthy, while stylistic or punctuation anomalies are easier to spot in side-by-side comparisons yet remain risky in real-world, zero-knowledge settings. This highlights the challenge of building robust defenses against diverse, evolving attack methods.
Appendix I Ablation Study – Impact of Aspect Tags in Prompts on ACOV Scores
Considering the ACOV advantage may stem from the structured prompt, in particular, explicit aspect tags may guide models to address each dimension systematically, thereby increasing coverage relative to human-written reviews. To assess the extent of this effect, we perform an ablation study in which aspect tags are either included or removed from the prompts. Results are reported in Table 10.
Table 10: Ablation study on the PeerRead dataset examining the impact of explicit aspect tags in prompts on ACOV scores. “w/ tag” and “w/o tag” denote prompts with and without explicit aspect tags, respectively. “Llama” refers to Llama-3.3-70B, and “Mistral” to Mistral-small-3.1.
Models GPT-4o-mini GPT-4o Llama Mistral
w/ tag 0.6066 0.9154 0.6489 0.9202
w/o tag 0.3556 0.4317 0.1702 0.4024
Across all models, the removal of explicit aspect tags leads to a decrease in ACOV. This indicates that structured prompt design plays a central role in achieving higher coverage. Consistent with prior work Yuan et al. (2022), we emphasize that ACOV should be interpreted as a coverage-oriented metric, not as a direct indicator of human-level review quality. The ablation highlights that, when explicitly guided, LLMs are able to systematically cover the intended review template, underscoring their potential as tools for generating structured and comprehensive feedback.
Appendix J Ablation Study – Impact of Attack Focus Localization
To evaluate the contribution of the proposed Attack Focus Localization algorithm, we conduct an ablation study that examines its role in improving adversarial effectiveness on long scientific documents. Unlike previous black-box adversarial methods that primarily target short inputs, the proposed approach identifies critical regions within lengthy academic papers and applies perturbations selectively. This localized strategy allows perturbations to be concentrated on semantically important sections, thereby enhancing attack effectiveness.
Table 11 presents the results of this ablation study on the PeerRead dataset. For each attack method, we compare the Attack Success Rate (ASR) and the Average Score Shift when perturbations are applied to the entire document without localization versus when using our proposed strategy.
Table 11: Ablation study on the impact of Attack Focus Localization on adversarial effectiveness. Values are presented as (Without/With Attack Focus Localization).
Attacks ASR $\uparrow$ Average Score Shift $\uparrow$
DeepWordBug 0.11 / 0.92 -0.12 / 4.27
PuncAttack 0.10 / 0.87 -0.11 / 3.90
TextFooler 0.10 / 0.92 0.22 / 4.61
BERT-Attack 0.10 / 0.85 0.02 / 4.13
StyleAdv 0.10 / 0.83 -0.10 / 3.58
The results show that incorporating the proposed Attack Focus Localization algorithm consistently boosts both ASR and Average Score Shift across all attack methods. These findings demonstrate that this localization strategy serves as a generalizable and effective approach for adversarial attacks on long-form academic documents, highlighting its importance in evaluating the robustness of LLM-based peer review systems.